Suppose Lauren, Leslie and Lydia all purchase tablets for R1 500 each. Lauren's willingness to pay was R3 500, Leslie's willingness to pay was R2 500 and Lydia's willingness to pay was R2 000. Total consumer surplus for these three would be


1
Expert's answer
2021-06-07T10:34:17-0400

If P is market price, then total consumer surplus is:

"CS = (350 - P) + (300 - P) + (250 - P) = 450."


"900 - 3P = 450"


"P = R150."